OpenStudy (anonymous):

[2 3] times [2 -7] [1 1] [3 -2] Those are supposed to be matrices can anyone help me solve this?

OpenStudy (anonymous):

like they are on top of each other beacause i dnt kno how to make a matix on here but they are 2x2 matrices

OpenStudy (anonymous):

let A=\[\left[\begin{matrix}2 & 3 \\ 1 & 1\end{matrix}\right]\] and B=\[\left[\begin{matrix}2 & -7 \\ 3 & -2\end{matrix}\right]\] then AB=\[[A(first column of B)|A(secondcolumn of B)]\] =\[\left[\begin{matrix}13 & -20 \\ 5 & -9\end{matrix}\right]\]

OpenStudy (anonymous):

how did u do those brackets?

OpenStudy (anonymous):

i dnt get how u got tht answer

OpenStudy (anonymous):

See, multiply the corresponding elements of two matrices given and tell me what you get after multiplying??

OpenStudy (anonymous):

dnt i do like 2x2 then 3x-7? like tht in order? because 2x2 is not 13 nd i dnt kno whea he got tht from

OpenStudy (anonymous):

See, I first give you my example then do this question on your own..  Getting or I should explain more??